QUESTION 12

In an economy in which real output grows at an average rate of 3 percent per year, a 7 percent average rate of growth in the money supply would result in a(n): a. $7 decrease in the price level each year. b. inflation rate of 4 percent, if the velocity of money in circulation is constant. c. increase in the velocity of money in circulation. d. inflation rate of -4 percent, if the velocity of money in circulation is constant. e. $7 increase in the price level each year.
Answer #1

(12) According to Quantity Theory of Money.

M * V = P * Y

Where, M is the money supply.

V is the velocity of money

P is the price level.

Y is the real GDP.

In terms of growth:

% change in M + % change in V = % change in P + % Change in Y

We assume that there is constant velocity of money. So the % change in V would be zero.

7 + 0 = % change in P + 3

% change in P = 7 - 3

% change in P = 4.

There would be an inflation rate of 4%, if the velocity of money in circulation is constant.

Answer: Option (B)